# 3DXTalker: Unifying Identity, Lip Sync, Emotion, and Spatial Dynamics in Expressive 3D Talking Avatars

### v2

@@ 4.3 Ablation Study
4.4 Analysis
Global Emotion Scalability. To further examine the model’s emotion controllability, we compute cosine similarity across seven emotions in the FLAME expression space, as listed in Table 3. 3DXTalker consistently achieves higher similarity scores than DEEPTalk, demonstrating superior emotion controllability and coherence. Figure 6 shows temporal emotion controllability. In Figure 6(a) Without explicit emotion conditioning control, the generated talking head stays near a neutral state with limited affective variation. In contrast, Figure 6(b) shows that our hybrid transition control enables smooth, flicker-free transitions across multiple emotion categories while preserving identity-consistent geometry, indicating that the control signal modulates expression in a structured and continuous manner. Moreover, we further present a t-SNE visualization of our expression predictions in Appendix  E. Partial overlaps between angry–disgust and surprise–fear are observed due to their similar facial activation patterns, while other clusters remain well separated, indicating that our model captures meaningful structure in the FLAME expression space.
Pose Analysis. We further analyze both jaw and head pose predictions. For jaw pose (Figure 7), the predicted jaw-opening pose ($\boldsymbol{\theta}_{0}^{j}$) trajectories with two individual seeds exhibit temporal consistence with the ground truth and closely follow the audio-amplitude, demonstrating that amplitude conditioning effectively contributes to mouth aperture. Regarding head pose (Figure 6), 3DXTalker maintains high stability and continuity across both supported modes. The results demonstrate that the model achieves both realistic pose generation and scalable, user-controllable dynamics without introducing abrupt jitter or pose discontinuities.
@@ 5 Conclusion
In summary, we present 3DXTalker, a comprehensive framework for expressive audio-driven 3D talking avatar generation that jointly advances identity modeling, audio expressiveness, and controllable dynamics. By constructing a scalable 2D-to-3D data-curated identity modeling pipeline with disentangled representations, 3DXTalker enables robust generalization across diverse speaker identities without relying on costly 3D capture. Moreover, by introducing audio-rich representations that explicitly capture frame-wise amplitude and emotional cues, our method achieves accurate lip synchronization and nuanced emotional expression. Combined with controllable head-pose modulation built upon natural motion priors, 3DXTalker produces identity-consistent, lip-synchronized, emotionally expressive, and spatially dynamic 3D talking avatars within a unified paradigm. Overall, our results demonstrate the effectiveness of the proposed formulation and highlight a practical path toward holistic expressivity in next-generation avatar systems. We believe that 3DXTalker provides a flexible and extensible foundation for future research on expressive digital humans, as well as broader applications in virtual communication and digital content creation.
Appendix A EMOCA Modeling Preliminary
@@
To leverage the abundant identities, emotional styles, and dynamic motion patterns available in large-scale 2D video data, we adopt the EMOCA parametric autoencoder [14], which lifts individual video frames into a controllable 3D facial parameter space, as illustrated in Figure 8. This formulation allows us to effectively bridge the gap between rich 2D visual observations and structured 3D facial representations, enabling scalable identity and motion modeling without relying on explicit 3D capture.
Specifically, the model encodes each 2D image into a set of FLAME parameters [27], including the identity-dependent shape $\boldsymbol{\beta}\in\mathbb{R}^{100}$, head pose $\boldsymbol{\theta}\in\mathbb{R}^{6}$, facial expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$, as well as an additional detail parameter $\boldsymbol{\delta}\in\mathbb{R}^{128}$ that captures high-frequency facial geometry. These parameters are then decoded into a coarse facial mesh $\mathbf{M}{\text{coa}}\in\mathbb{R}^{5023\times 3}$ by deforming the FLAME template $\mathbf{\bar{V}}$ according to $\boldsymbol{\beta}$, $\boldsymbol{\psi}$, and $\boldsymbol{\theta}$. Finally, the coarse mesh is further refined by a detail decoder $\mathcal{D}{\text{det}}$}, which applies a facial displacement map to generate a high-resolution mesh $\mathbf{M}_{\text{det}}\in\mathbb{R}^{59315\times 3}$, formally defined as:
where $\mathbf{V}_{\text{int}}$ denotes an intermediate mesh representation, $\text{B}_{S}$ and $\text{B}E$ are the shaping and expression blend functions, respectively, and LBS refers to the linear blend skinning operation parameterized by the joint regressors $\text{J}P$ and the skinning weights $\mathcal{W}$. The template UV map is represented by $\mathbf{\bar{U}}$, and Fdet applies the predicted facial displacement map to the coarse mesh $\mathbf{M}{\text{coa}}$, refining it into a detailed surface geometry. Notably, the pose parameter $\boldsymbol{\theta}$ can be naturally decomposed into head pose and jaw pose, enabling head motion to be modeled and manipulated independently from facial articulation through simple linear variations in the pose space. This separation facilitates explicit and interpretable control over global head dynamics without interfering with lip and expression movements. Overall, this parameterization yields a disentangled and physically meaningful 3D facial representation, which forms the foundation of our subsequent data-curated identity modeling pipeline and supports scalable learning across diverse identities, expressions, and motion patterns.
@@ Appendix B Implementation Details
B.1 Dataset Curation Pipeline Details
To address the scarcity of training data and limited identity diversity in 3D talking avatar generation, we construct a large-scale corpus by integrating six diverse 2D video datasets. Specifically, we collect six widely used 2D talking video datasets, including three lab-controlled datasets (GRID [11], RAVDESS [29], MEAD [46]) and three in-the-wild datasets (VoxCeleb2 [10], HDTF [52], CelebV-HQ [53]). The lab-controlled datasets provide high-quality recordings with articulated facial movements and emotional expressions, while the in-the-wild datasets introduce a broad range of identities, speaking styles, and natural head pose dynamics. To ensure consistency and quality across these sources, we apply a unified data preprocessing pipeline to filter outliers, following the steps below:
Table 4: Statistics of the curated talking video dataset after our data curation pipeline.
@@ B.1 Dataset Curation Pipeline Details
  1. (1)
  2. (2)
    Signal-to-Noise Ratio Filtering. To remove clips compromised by strong background noise, music, or environmental interference, we compute the signal-to-noise ratio (SNR) for each audio segment, discarding samples with SNR below a predefined threshold. This step is critical for in-the-wild datasets, where recordings often contain crowd noise, reverberation, or microphone artifacts. The SNR filtering of speech signals ensures reliable cues for reliable amplitude extraction, emotion inference, and audio–visual synchronization.
  3. (3)
    Language Filtering. We enforce linguistic consistency by filtering clips based on spoken language using Whisper [37], discarding non-English samples or those with low detection confidence. This filtering step prevents mixing heterogeneous phonetic structures and prosodic patterns, reducing potential interference with audio–visual alignment.
  4. (4)
    Audio-Visual Sync Filtering. To guarantee strict audio–visual alignment, we filter samples using SyncNet [9] to evaluate the temporal correlation between lip motion and the speech signal. We discard clips exhibiting high synchronization errors, as well as those containing abrupt scene cuts or off-screen speakers (e.g., voice-overs). Eliminating these misaligned pairs is essential, as they provide incorrect supervision signals that can significantly degrade lip-sync learning and hinder expression modeling.
  5. (5)
@@ B.2 Setup Details
The Transformer backbone of 3DXTalker is designed to balance modeling capacity and computational efficiency. It comprises 6 Transformer blocks with a hidden dimension of 768, enabling effective integration of identity, audio, and temporal cues. To better disentangle downstream targets, each prediction head (identity, pose, and expression) is equipped with two additional Transformer blocks that function as lightweight, task-specialized decoders. This hierarchical architecture ensures that shared representations capture global consistency, while branch-specific modules refine modality-dependent details.
For training, 3DXTalker is optimized for 100 epochs using 250-frame sequences and a batch size of 128, which provides a sufficient temporal window for modeling natural motions. We adopt AdamW with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.01, coupled with a OneCycleLR schedule to stabilize warm-up and improve convergence. The flow-matching module operates with 512 steps during training to ensure high-fidelity trajectory learning, while 32 inference steps are used at test time to achieve a favorable balance between accuracy and generation speed.
Experiments are conducted on a shared high-performance computing cluster equipped with NVIDIA H100 GPUs. The complete training of 3DXTalker utilizes two H100 GPUs with BF16 mixed precision, resulting in a total training time of approximately five hours. Despite the shared-cluster variability, this setup demonstrates that 3DXTalker is computationally efficient and scalable for large-scale expressive talking avatar generation.
@@ B.3 Metrics Details
  • Mean Rank (MR) is calculated by averaging the rankings obtained from a subjective user study. As shown in Figure 9, we designed an interactive interface where participants were presented with a ground-truth identity and eight anonymized, shuffled videos generated by the comparative models. Evaluators are instructed to watch all videos and rank them from 1 (Best) to 8 (Worst) based on three criteria: identity consistency, lip synchronization, and emotional expression.
Appendix C Baseline Visualizations
@@ 2. For each session $s$ containing emotion $e$:
Figure 11: t-SNE visualization of our predicted expression. Partial overlaps between angry–disgust and surprise–fear correspond to their naturally similar facial activation patterns.
Appendix E Emotion Expression
E.1 Emotion Expression Template
To enable explicit global emotion control in addition to audio-driven fine-grained dynamics, we derive emotion templates directly from the FLAME expression parameter subspace. These templates represent canonical expression directions for seven basic emotions and are used for emotion scaling and interpolation during inference. MEAD dataset provides videos recorded under controlled lighting and contains high-intensity emotion performances, making it suitable for learning expression prototypes. Inspired by this, we obtain expression templates by analyzing the FLAME expression parameters ($\boldsymbol{\psi}\in\mathbb{R}^{50}$) extracted by EMOCA from the MEAD dataset, with the full procedure detailed in Table 6. The extracted expression templates cover seven typical emotions (Angry, Contempt, Disgust, Fear, Happy, Sad, and Surprise). We further introduce a global scaling factor $\alpha\in\{1.0,1.2,1.4,1.6,1.8,2.0\}$ to control emotion intensity, visualized in Figure 12. Naturally, some emotions (e.g., angry–disgust and surprise–fear) share similar facial muscle activation patterns, which explains the observed partial overlaps in the t-SNE space, whereas the remaining emotions form well-separated clusters (present in Figure 11). During inference, we can adjust the global emotional tone by interpolating between the reference expression $\boldsymbol{\psi}_{ref}$ and the scaled emotion template:
@@ E.1 Emotion Expression Template
Figure 13: Additional qualitative comparisons of four emotion categories (Disgust, Contempt, Fear, Surprise) across representative baselines. 3DXTalker generates more expressive emotion patterns with clearer facial activations than DEEPTalk and EMOTE.
@@
E.2 More Emotion Comparisons
To further demonstrate our model’s emotion expressivity, we present additional qualitative comparisons across four representative emotion categories, as shown in Figure 13. Across all categories, our model produces more expressive and coherent facial deformations, demonstrating improved emotion disentanglement and modulation compared with DEEPTalk [25] and EMOTE [15].
Appendix F Head Pose Dynamics
[…]
Refer to caption Refer to caption
Appendix G Downstream Application
We employ ComfyUI to achieve texture mapping on the talking avatars generated by our model, using the Wan 2.2 diffusion model [45]. Wan is a versatile video generation framework that supports text-to-video, image-to-video, and video-to-video synthesis. To align with 3DXTalker’s needs, we adopt two variants, Wan2.2-Fun-Control and Wan2.2-Speech-to-Video (S2V). Fun-Control enables fine-grained video control through depth, pose, and edge guidance, complemented by LLM-generated prompts. Specifically, we render a depth video from our 3D mesh sequence and feed it into Fun-Control to drive head pose, lip movement, and emotion-consistent facial dynamics. Meanwhile, the reference image provides identity-appearance cues, ensuring that the synthesized video maintains the subject’s visual characteristics, while the input audio provides synchronized speech. In contrast, S2V is an audio-driven video generation model that conditions on text prompts, reference images, and speech signals. It specializes in transforming a static portrait and audio into a synchronized talking video. However, unlike Fun-Control, S2V does not utilize depth guidance, which makes it less reliable in producing accurate head-pose dynamics and temporally coherent motion.
Figure 16 presents the rendered results produced using the Wan model. For each video, Wan requires two types of prompts: a positive prompt, which directs the model toward desired visual and motion characteristics, and a negative prompt, which constrains unwanted artifacts or behaviors during generation. In our setup, we vary only the positive prompts to tailor each video to its intended scene or stylistic effect while keeping the default negative prompt unchanged. This setup follows Wan’s recommended usage and ensures consistent, stable video generation across different rendering configurations. The positive prompts used for Wan rendering in Figure 16 are provided below:
@@ Appendix G Downstream Application
Appendix H Prompt Design for Motion Trajectory Control of Head Pose
To enrich spatial dynamics beyond the subtle audio-driven motion produced by 3DXTalker, we introduce an LLM-driven Cinematography strategy. This module controls head-pose motion and camera movement separately by taking an audio clip and a user prompt as input, returning smooth, interpretable motion functions that can be applied to mesh-sequence rendering. The key idea is to leverage the reasoning and generative capabilities of large language models to convert high-level textual descriptions, such as “energetic presentation”, “subtle and calm”, or “cinematic orbit shot”, into executable motion trajectories. By analyzing both the acoustic rhythm and the semantic intent provided in the prompt, the module produces parameterized functions that modulate head orientation or camera position over time, enabling expressive speaking styles and professional cinematographic effects without hand-crafted animation. This design significantly expands the range of spatial behaviors achievable by 3DXTalker, offering fine-grained control and diverse visual dynamics while preserving coherence with the underlying mesh animation. The prompt templates for both head-pose and camera-motion control are provided below.
Appendix I Discussion
While 3DXTalker demonstrates strong performance in expressive audio-driven 3D talking avatar generation, several directions remain open for future improvement.
First, the quality of the generated facial motion is inherently influenced by the modeling capacity of the underlying parametric face representation. In this work, we rely on EMOCA to lift 2D video frames into the FLAME parameter space, which enables scalable identity and motion modeling. However, the expressiveness and fidelity of the final results are ultimately bounded by the representational power of EMOCA and FLAME, particularly for subtle facial details and extreme expressions. Advances in parametric face modeling or hybrid representations may further enhance the realism and expressivity of the generated avatars.
Second, although our framework focuses on audio-driven generation, human expressive behavior is not solely determined by speech signals. Additional non-audio cues, such as linguistic content, discourse structure, or higher-level communicative intent, also play an important role in shaping expressive delivery. Incorporating such complementary signals may further enrich expressive diversity and improve alignment with user intent.
Finally, while we introduce audio-rich representations that capture frame-wise amplitude and emotion cues, modeling fine-grained emotional dynamics solely from speech remains a challenging problem. Emotional expression in speech is often subtle, context-dependent, and temporally evolving, which may not be fully captured by current audio features alone. Future work could explore more advanced affect modeling or multi-scale emotion representations to better reflect the complexity of speech-driven emotions.
Figure 17: Visualization comparisons illustrating how mouth-aperture patterns align with phonetic symbols across different models.

### v3

@@ 5 Conclusion
In this paper, we present 3DXTalker, an integrated flow-matching framework for expressive audio-driven 3D talking avatar generation. By combining scalable 2D-to-3D data integration, audio-rich representations, and plug-in semantic controllability, 3DXTalker jointly improves identity consistency, lip synchronization, emotional expression, and head-pose dynamics within a unified framework. This design enables expressive 3D talking avatars to better balance subject fidelity, speech alignment, emotional expressiveness, and controllable spatial dynamics in a coordinated manner. Extensive experiments demonstrate that 3DXTalker achieves superior and more comprehensive performance in expressive 3D talking avatar generation, validating both its effectiveness and controllability. Overall, our work takes a practical step toward holistic expressivity in next-generation 3D avatar systems and provides a flexible foundation for future research on expressive digital humans.
Appendix A EMOCA Preliminary
@@
To leverage the abundant identities, emotional styles, and motion patterns available in large-scale 2D video data, we adopt EMOCA [13] as a parametric 2D-to-3D facial reconstruction model. As illustrated in Figure 10, EMOCA lifts each video frame into a structured FLAME parameter space, including shape, expression, pose, and detail codes, and further decodes them into a detailed 3D facial mesh. This formulation provides a controllable bridge from rich 2D visual observations to structured 3D facial representations, enabling scalable identity-aware modeling without requiring explicit 3D motion capture. It also offers a disentangled parameterization that is well suited for subsequent temporal motion modeling and controllable 3D avatar generation.
Table 4: Comparison of existing 3D talking-head datasets and our dataset. Our dataset provides the most comprehensive 3D fitting FLAME annotations, including identity, expression, pose, and detail, while also covering both lab and wild scenarios with speech and song audio. BS denotes blendshape coefficients, and ARKit refers to Apple’s facial motion capture parameterization.
Dataset Representation Dimension Annotation Source Audio Sequence Duration FPS Subject Identity Expression Pose Detail
[…]
BEAT [30] BS 52 ARKit Lab Speech 2,508 76h 60 30 - - - -
Song2face [25] BS 51 ARKit Lab Song - 1.93h - 7 - - - -
3D-ETF [37] BS 52 3D fitting Lab & Wild Speech 3,479 6.97h 30 165 - - - -
MMHead [52] FLAME 56 3D fitting Wild Speech 35,903 49h 25 2k+
TFHP [45] FLAME 56 3D fitting Wild Speech 1,052 26.5h 25 588
Ours FLAME 284 3D fitting Lab & Wild Speech&Song 11,706 67.41h 25 2k+
@@ Appendix B Implementation Details
B.1 Scalable Data Integration Pipeline
To address the scarcity of training data and limited identity diversity in 3D talking avatar generation, we construct a large-scale corpus by integrating six diverse 2D video datasets. Specifically, we collect six widely used 2D talking video datasets, including three lab-controlled datasets (GRID [11], RAVDESS [31], MEAD [49]) and three in-the-wild datasets (VoxCeleb2 [10], HDTF [56], CelebV-HQ [57]). The lab-controlled datasets provide high-quality recordings with articulated facial movements and emotional expressions, while the in-the-wild datasets introduce a broad range of identities, speaking styles, and natural head pose dynamics. A comprehensive comparison of our dataset with existing 3D talking-head datasets is presented in Table 4. Compared with prior vertex-based and blendshape-based datasets, our dataset is built in the FLAME parameter space, providing a more compact and semantically structured representation for controllable 3D facial animation. Moreover, among existing FLAME-based datasets, our dataset offers the most comprehensive annotations, jointly covering identity, expression, pose, and detail. It also maintains strong diversity in data sources, audio types, and subject coverage, making it a more suitable benchmark for expressive 3D talking avatar generation.
Table 5: Statistics of the constructed 3D talking-head dataset via our data integration pipeline.
@@ B.1 Scalable Data Integration Pipeline
To ensure consistency and quality across talking video sources, we apply a unified data preprocessing pipeline to filter outliers, following main steps below:
  1. (1)
  2. (2)
    Signal-to-Noise Ratio Filtering. To remove clips compromised by strong background noise, music, or environmental interference, we compute the signal-to-noise ratio (SNR) for each audio segment, discarding samples with SNR below a predefined threshold. This step is critical for in-the-wild datasets, where recordings often contain crowd noise, reverberation, or microphone artifacts. The SNR filtering of speech signals ensures reliable cues for reliable amplitude extraction, emotion inference, and audio–visual synchronization. By suppressing acoustically corrupted samples at the preprocessing stage, we further improve the robustness and consistency of the downstream speech-driven facial motion generation pipeline.
  3. (3)
    Language Filtering. We enforce linguistic consistency by filtering clips based on spoken language using Whisper [39], discarding non-English samples or those with low detection confidence. We retain only English videos to reduce cross-lingual variation in phonetic structure and prosodic patterns, which would otherwise introduce additional difficulty for learning stable audio-visual alignment. This step helps provide more consistent supervision for lip synchronization, emotion modeling, and speech-driven facial dynamics. This choice is made to improve training consistency rather than to restrict the framework to English-only applications. Extending to multi-lingual datasets is expected to further improve generalization and enable broader applicability.
  4. (4)
    Audio-Visual Sync Filtering. To guarantee strict audio to visual alignment, we filter samples using SyncNet [9] to evaluate the temporal correlation between lip motion and the speech signal. We discard clips exhibiting high synchronization errors, as well as those containing abrupt scene cuts or off-screen speakers (e.g., voice-overs). Eliminating these misaligned pairs is essential, as they provide incorrect supervision signals that can significantly degrade lip-sync learning and hinder expression modeling.
  5. (5)
@@ B.2 Setup Details
We use EMOCA V2 to lift 2D videos into FLAME representations, with the pretrained weight EMOCA_v2_lr_mse_20. For audio encoding, we adopt the pretrained models wavlm-base-plus and emotion2vec_plu_base to extract global audio embeddings and frame-wise emotion features from input audio, respectively. These modules provide the structured 3D facial parameters and audio-rich cues used by the generative backbone.
The integrated audio-rich backbone of 3DXTalker consists of 6 diffusion transformer blocks with a hidden size of 768. Each prediction head is implemented with 2 additional blocks to refine branch-specific outputs from the shared hidden representation. In this way, the shared backbone models global identity, audio, and temporal dependencies, while the prediction heads specialize in different output factors.
Experiments are conducted on two NVIDIA H100 GPUs. For training, 3DXTalker is optimized for 100 epochs using 250-frame sequences and a batch size of 128. We adopt AdamW with a learning rate of $1\times 10^{-4}$ and a weight decay of 0.01, together with a OneCycleLR scheduler. The flow-matching module uses 512 steps during training and 32 steps during inference, which provides a favorable balance between trajectory fidelity and computational efficiency. Following prior works [37, 45], we further apply a Savitzky–Golay filter to the predicted sequences during post-processing to improve motion smoothness.
During training, the sequence length is set to 250 frames. At inference, longer sequences are generated via a sliding-window strategy. Because each window is defined relative to a shared reference frame, identity is implicitly recalibrated across windows. In our experiments, outputs remain stable without noticeable degradation for sequences up to 10 minutes under typical talking-head settings. Longer durations are likely feasible, but were not further explored due to computational cost, indicating a soft rather than strict practical limit on sequence length.
@@ B.3 Metrics Details
  • Mean Rank (MR) is calculated by averaging the rankings obtained from a subjective user study. As shown in Figure 11, we designed an interactive interface where participants were presented with a ground-truth identity and eight anonymized, shuffled videos generated by the comparative models. Evaluators are instructed to assess all videos and rank them from 1 (Best) to 8 (Worst) based on three criteria: identity consistency, lip synchronization, and emotional expression.
Figure 12: Visualization comparisons illustrating how mouth-aperture patterns align with phonetic symbols across models.
Appendix C Full Baseline Comparisons
@@ 2. For each session $s$ containing emotion $e$:
Appendix D Emotion Expression
D.1 Semantic Emotion Control
To enable explicit global emotion control in addition to audio-driven fine-grained dynamics, we derive emotion templates directly from the FLAME expression parameter subspace. These templates represent canonical expression directions for seven basic emotions and are used for emotion scaling and interpolation during inference. Furthermore, MEAD dataset provides videos recorded under controlled lighting and contains high-intensity emotion performances, making it suitable for learning expression prototypes. Inspired by this, we obtain expression templates by analyzing the FLAME expression parameters ($\boldsymbol{\psi}\in\mathbb{R}^{50}$) extracted by EMOCA from the MEAD dataset, with the full procedure detailed in Table 7. The extracted expression templates cover seven typical emotions (Angry, Contempt, Disgust, Fear, Happy, Sad, and Surprise). We further introduce a global scaling factor $\alpha\in\{1.0,1.2,1.4,1.6,1.8,2.0\}$ to control emotion intensity, visualized in Figure 13. During inference, we can adjust the global emotional tone by interpolating between the reference expression $\boldsymbol{\psi}_{ref}$ and the scaled emotion template:
@@ D.1 Semantic Emotion Control
D.2 More Emotion Visualization Comparisons
To further demonstrate our model’s emotion expressivity, we further present qualitative comparisons across additional four representative emotion categories, as shown in Figure 14. Across all categories, our model produces more expressive and coherent facial deformations, demonstrating improved emotion disentanglement and modulation compared with DEEPTalk [27] and EMOTE [14]. In particular, for more intense emotions such as Disgust and Fear, our model produces stronger deformations around key facial regions, including the mouth, eyes, and brows. For Contempt and Surprise, it also maintains better coordination among local facial movements, resulting in more coherent overall expressions. In comparison, DEEPTalk and EMOTE often generate weaker or less distinctive emotional patterns. These observations further support that 3DXTalker provides stronger emotion disentanglement and more effective emotion modulation in expressive 3D talking avatar generation. These comparisons suggest that our framework better preserves category-specific emotional characteristics while maintaining facial coherence. This improvement is consistent with our design choice of combining disentangled FLAME-based modeling with frame-wise emotion-aware conditioning, which helps separate emotional modulation from identity and other motion factors.
Figure 14: More qualitative comparisons of additional four emotion categories (Disgust, Contempt, Fear, Surprise) across representative baselines. 3DXTalker generates more expressive emotion patterns with clearer facial activations than DEEPTalk and EMOTE.
@@ D.2 More Emotion Visualization Comparisons
Appendix E Head Pose Dynamics
E.1 Semantic Pose Control
To complement emotion semantic control, we further introduce semantic control over head-pose dynamics through an LLM-driven prompting strategy. Given the driving audio and a text prompt describing the desired presentation style, the language model generates a smooth and interpretable head-pose function for FLAME global pose parameters. The goal is to convert high-level semantic descriptions, such as “energetic presentation”, “subtle and calm”, or “stage-presenting delivery”, into executable head-pose trajectories. The prompt template used for this process is provided below. By combining acoustic cues from the input waveform with the semantic intent specified in the prompt, the generated trajectory modulates head orientation over time in a smooth and controllable manner. Rather than replacing the model-learned natural head motion, the generated trajectory is superimposed onto it, thereby preserving realism while enabling semantically meaningful stylistic variation. As a result, 3DXTalker supports semantically meaningful head-pose variation while maintaining identity consistency and expression dynamics.
@@ E.1 Semantic Pose Control
E.2 More Head-Pose Visualization Comparisons
[…]
Appendix F More Amplitude Analysis Cases
Refer to caption Refer to caption
Appendix G Downstream Application: Wan 2.2
We employ ComfyUI to achieve texture mapping on the talking avatars generated by our model, using the Wan 2.2 [48]. Wan 2.2 is a versatile video generation framework that supports text-to-video, image-to-video, and video-to-video synthesis. To align with 3DXTalker’s needs, we adopt two variants, Wan2.2-Fun-Control and Wan2.2-Speech-to-Video (S2V). Fun-Control enables fine-grained video control through depth, pose, and edge guidance, complemented by LLM-generated prompts. Specifically, we render a depth video from our 3D mesh sequence and feed it into Fun-Control to drive head pose, lip movement, and emotion-consistent facial dynamics. Meanwhile, the reference image provides identity cues, ensuring that the synthesized video maintains the subject’s visual characteristics, while the input audio provides synchronized speech. In contrast, S2V is an audio-driven video generation model that conditions on text prompts, reference images, and speech signals. It specializes in transforming a static portrait and audio into a synchronized talking video. However, unlike Fun-Control, S2V does not utilize depth guidance, which makes it less reliable in producing accurate head-pose dynamics and temporally coherent motion.
Figure 18 presents the rendered results produced using the Wan 2.2 model. Wan 2.2 requires two types of prompts: a positive prompt, which directs the model toward desired visual and motion characteristics, and a negative prompt, which constrains unwanted artifacts or behaviors during generation. In our setup, we vary only the positive prompts to tailor each video to its intended scene or stylistic effect while keeping the default negative prompt unchanged. For the full-control setting, the positive prompts are used together with the rendered depth videos and reference images to encourage identity-consistent and visually coherent talking video synthesis. The positive prompts for Figure 18 are presented below:
Appendix H Future Work
Discussion on EMOCA Modeling. As the upstream 2D-to-3D reconstruction module, EMOCA may introduce errors under challenging conditions such as occlusion, motion blur, extreme poses, or low-resolution facial regions. These inaccuracies can affect the recovered FLAME parameters, including shape, expression, pose, and detail, and may propagate to downstream motion generation. In particular, identity-related estimation errors may weaken identity consistency, while noisy expression or pose parameters may reduce motion stability. Despite these limitations, EMOCA currently provides one of the most comprehensive monocular pipelines for jointly estimating identity, expression, pose, and detail parameters in a unified framework, making it well-suited for our task setting. While our pipeline incorporates standard data filtering and temporal smoothing steps that can partially reduce high-frequency artifacts, the reconstruction accuracy remains fundamentally constrained by the upstream lifting process. Improving the robustness of 2D-to-3D reconstruction, especially under challenging visual conditions, represents an important direction for future work.
Discussion on Dataset. Although our training dataset is already relatively large-scale, the achievable level of fine-grained expressivity remains influenced by the diversity and quality of available data. Subtle speech-dependent facial dynamics, such as nuanced articulation patterns and culturally influenced expression styles, require sufficiently rich supervision in both audio and parameter spaces. Expanding the dataset with more fine-grained motion variations and broader linguistic diversity, particularly non-English speech data, could further improve the model’s ability to capture subtle expression differences and generalize across identities and speaking styles. Therefore, the upper bound of expressive controllability is determined not only by the representation capacity but also by the diversity and coverage of training data. Exploring richer multi-lingual and fine-grained motion datasets remains an important direction for future work to further enhance expressive fidelity.